OpenStudy (anonymous):

If f(X) = 25 -x ^2 and g(x) = x+5, what is (f/g) (x)?

OpenStudy (anonymous):

help please i dont understand

jhonyy9 (jhonyy9):

so f/g (x) what mean ? f(x) divide by g(x) yes ?

OpenStudy (anonymous):

I dont understand?

jhonyy9 (jhonyy9):

so f(x) = 25-x^2 then you know formula a^2 -b^2 =(a-b)(a+b) so you can using it here because you can rewriting f(x) = (5-x)(x+5) ok ?

jhonyy9 (jhonyy9):

and now than you rewrite f/g(x) how will be ?so using this factored form of f(x)

OpenStudy (anonymous):

5x+25-x^2-5x??

OpenStudy (anonymous):

yeah i still dont get what the answer is ??

jhonyy9 (jhonyy9):

so will be (5-x)(5+x) ---------- = ? x+5

jhonyy9 (jhonyy9):

now this is easy just simplifie what you can ok ?

OpenStudy (anonymous):

x+5?? is that the answer

jhonyy9 (jhonyy9):

why x+5 ? when there is down and up too ? this will be simplified so then what will remained there ?

OpenStudy (anonymous):

i give up

jhonyy9 (jhonyy9):

(5-x)(5+x) ---------- than simplifie 5+x by 5+x what will get ? 5+x

jhonyy9 (jhonyy9):

5+x from numerator will be simplified by 5+x from denominator so than will get 5-x yes ?

jhonyy9 (jhonyy9):

hope you understand it now good luck bye